RC Meeting, Thursday, September 30 , 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 am, remote

  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Accelerator Status

  1. Sep 29 Swing: Accelerator first attempt to deliver beam to C was delayed
  2. Sep 30 Owl: Accelerator delivered ~45 min. of useful beam at 30 uA, mainly due to BLM trips
  3. Sep 30 Day: From 7:45 am meeting, Accelerator Recovery is expected to continue throughout the Day Shift and Beam Delivery is unlikely.

Hall C Status

  1. Sep 29 Swing: Harp Scans and Carbon Hole Check completed towards the end of Sep 29 Swing Shift
  2. Sep 30 Owl: From the ~45 min. of useful beam at 30 uA, ~18 pions were collected (Thanks to Jacob Murphy for providing info)

Current Run Plan

  1. Sep 30 Owl/Day/Swing (today):
    • Contine data-taking at current PION-LT kinematic setting ( low eps, Q2=6 GeV2, W=3.14, x=0.39 )
      • Data collected as of Sep 29 Day Shift: 4155 pions (41% of the statistics and 67% of the PAC approved beam charge accumulated)
      • After 85% of beam charge has been accumulated ( this is subject to change given the circumstances ), request to lower beam current to 45 uA to take the remaining data
        until Friday Day Shift --> https://logbooks.jlab.org/entry/3914672 ( PENDING )
  2. SHMS Rotation to 5.5 deg is contingent upon beam delivery to Hall C (time/date is still "to be determined" TBD)
    • Will contact Amy Comer and Radcon ahead of time (today) to co-ordinate entrance to Hall ( PENDING )
      • Rotation will require several 3 spotters (1 in the pivot + 2 on the floor between each spectrometer and beamline)
      • If possible, do rotation in "Controlled Access" (spotter in pivot will have to use ALARA to minimize time near radiation area), otherwise, will go into "Restricted Access" (takes a bit longer)

Current Issues

  1. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.
  2. Why did target draw so much coolant (30 grams / sec) and will this be a problem once Hall A starts running next week? ( CLEARED )
  3. HCDESK1 computer dead ( CLEARED )

Special Requests

  1. Contact Ellen if there is an opportunity to access the hall (longer than 4 hours). This is for diagnosing the SHMS power supply.